OpenStudy (anonymous):

Will give medal!What is the resulting ordered pair if the value of the independent variable is 2? f(x) = –5x + 4

terenzreignz (terenzreignz):

if you have f(x), then the independent variable is x. So if x = 2, what is f(x)? \[\Large f(2) =\color{red}?\]

OpenStudy (anonymous):

2f..?

terenzreignz (terenzreignz):

You mustn't mistake f(x) as any sort of multiplication, no... f(x) = -5x + 4 f(a) simply means you replace the x with a. f(a) = -5a + 4 f(5) means you replace the x with 5... f(5) = -5(5) + 4 = -25 + 4 = -21 That said, what is f(2)= ?

OpenStudy (anonymous):

2?

terenzreignz (terenzreignz):

NO... like I said, replace the x in -5x + 4 with a 2, and what do you get?

OpenStudy (anonymous):

idk

terenzreignz (terenzreignz):

\[\Large f(\color{red}x) = -5\color{red}x +4\]\[\Large f(\color{red}2) = -5(\color{red}2)+4\] now solve :)

OpenStudy (anonymous):

14..?

OpenStudy (anonymous):

-14!

terenzreignz (terenzreignz):

It's a +4 there, not a -4 :P